# Demystifying LLM-as-a-Judge: Analytically Tractable Model for Inference-Time Scaling

### v1

@@ 4 Qualitative agreement with large language model reasoning
5 Conclusion
We introduce a simple, solvable model to study how inference-time scaling works, where multiple candidate answers are generated and chosen using a reward-based selection process. Our analysis shows that when the reward model is well aligned with the task, increasing the number of inference time samples reliably improves generalization, but when it is misaligned, there is an optimal finite number of samples and an optimal temperature for selection. In the best-of-k setting, we identify when extra inference-time compute is more valuable than additional training. Finally, our experiments with modern language models on a math problem benchmark validate theoretical results.
@@ A.1 Limit Laws for Maxima
$\mathbb{P}\!\left(\frac{M_{n}-b_{n}}{a_{n}}\leq x\right)\xrightarrow[n\to\infty]{}H(x)\qquad(x\in\mathbb{R}\text{ continuity points of }H),$
@@ A.2 Maximum Domains of Attraction and Norming Constants
$\lim_{n\to\infty}F^{n}(a_{n}x+b_{n})=H(x)\qquad(x\in\mathbb{R}\text{ continuity points of }H).$
@@
$\lim_{n\to\infty}n\bigl(1-F(a_{n}x+b_{n})\bigr)=-\ln H(x)\qquad(x\in\mathbb{R}).$
@@ Appendix B Proof of Result 1
$\delta=\mathbb{E}_{\mathbf{x}\sim\mathcal{N}(0,\bm{\Sigma})}\,\mathbb{E}_{y_{i}\sim\mathcal{N}(m(\mathbf{x}),s(\mathbf{x})^{2}),i=1,2,\dots,k}\left[\frac{\sum_{i=1}^{k}\big(y_{i}-\mu_{T}(\mathbf{x})\big)^{2}\,e^{-\big(y_{i}-\mu_{R}(\mathbf{x})\big)^{2}/T}}{\sum_{j=1}^{k}e^{-\big(y_{j}-\mu_{R}(\mathbf{x})\big)^{2}/T}}\right],$ (29)
to the leading order in $\sigma$. Here the posterior predictive has mean $m(\mathbf{x})$ and variance $\Sigma(\mathbf{x})=s(\mathbf{x})^{2}$ as follows
@@
$\hat{R}=R\Big(1-\alpha\,m_{\bm{\Sigma}}(R)\Big)=\frac{\sigma^{2}}{\gamma^{2}}\alpha,\qquad m_{\bm{\Sigma}}(R):=\frac{1}{d}\text{Tr}\!\big[\bm{\Sigma}(\bm{\Sigma}+RI)^{-1}\big].$ (32)
@@
$m(\mathbf{x})$ $=\bm{\mu}^{\top}\frac{\mathbf{x}}{\sqrt{d}}=\frac{1}{\sigma^{2}}\,\frac{\mathbf{x}}{\sqrt{d}}^{\top}\bm{\Omega}\sum_{i=1}^{n}y^{i}\frac{\mathbf{x}^{i}}{\sqrt{d}}=\frac{1}{\sigma^{2}}\,\frac{\mathbf{x}}{\sqrt{d}}^{\top}\bm{\Omega}\Big(\sum_{i=1}^{n}\frac{\mathbf{x}^{i}}{\sqrt{d}}\frac{\mathbf{x}^{i}}{\sqrt{d}}^{\!\top}\Big)\mathbf{w}_{T}\ +\ \frac{1}{\sigma^{2}}\,\frac{\mathbf{x}}{\sqrt{d}}^{\top}\bm{\Omega}\sum_{i=1}^{n}\eta^{i}\frac{\mathbf{x}^{i}}{\sqrt{d}}$
@@
$\mathrm{Var}\big(Z(\mathbf{x})\big)\;\xrightarrow[d,n\to\infty]{}\;\sigma^{2}\,\frac{\alpha\,m_{\bm{\Sigma}}^{(2)}(R)}{1-\alpha\,m_{\bm{\Sigma}}^{(2)}(R)},\quad m_{\bm{\Sigma}}^{(2)}(R):=\frac{1}{d}\text{Tr}\!\big[\bm{\Sigma}^{2}(\bm{\Sigma}+R\mathbf{I})^{-2}\big]$ (40)
@@
$\sum_{i=1}^{n}\frac{\mathbf{x}^{i}}{\sqrt{d}}{\frac{\mathbf{x}^{i}}{\sqrt{d}}}^{\top}=\frac{1}{d}\sum_{i=1}^{n}\mathbf{x}^{i}{\mathbf{x}^{i}}^{\top}=\frac{n}{d}\,\widehat{\bm{\Sigma}}=\frac{1}{\alpha}\,\widehat{\bm{\Sigma}},\quad\bm{\Omega}=\sigma^{2}\alpha(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}.$
@@
$\mathrm{Var}\big(Z(\mathbf{x})\mid\mathcal{D},\mathbf{x}\big)$ $=\frac{1}{\sigma^{2}\alpha}\,\frac{\mathbf{x}}{\sqrt{d}}^{\top}\bigl[\sigma^{2}\alpha(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\bigr]\widehat{\bm{\Sigma}}\bigl[\sigma^{2}\alpha(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\bigr]\frac{\mathbf{x}}{\sqrt{d}}$
@@
$\mathbb{E}_{\mathbf{x}}\big[\mathrm{Var}(Z(\mathbf{x})\mid\mathcal{D},\mathbf{x})\big]$ $=\sigma^{2}\alpha\,\frac{1}{d}\,\text{Tr}\Big((\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\bm{\Sigma}\Big).$ (43)
@@
$\mathrm{Var}\big(Z(\mathbf{x})\big)=\mathbb{E}_{\mathcal{D},\mathbf{x}}\big[\mathrm{Var}(Z(\mathbf{x})\mid\mathcal{D},\mathbf{x})\big],$
@@
$\mathrm{Var}\big(Z(\mathbf{x})\big)=\sigma^{2}\alpha\,\frac{1}{d}\,\text{Tr}\Big((\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\bm{\Sigma}\Big)\;\xrightarrow[d,n\to\infty]{}\;\sigma^{2}\,\frac{\alpha\,m_{\bm{\Sigma}}^{(2)}(R)}{1-\alpha\,m_{\bm{\Sigma}}^{(2)}(R)}.$
@@ Appendix C Proof of Result 2
$\delta\;=\;\mathbb{E}_{\mathbf{x}}\Big[\Delta_{T}(\mathbf{x})^{2}+s^{2}(\mathbf{x})+\Sigma_{l=1}^{3}(-1)^{l}\frac{C_{l}(\mathbf{x})}{t(\mathbf{x})^{l}}\prod_{i=1}^{l}\Big(1-\tfrac{i}{k}\Big)+\mathcal{O}\!\Big(t(\mathbf{x})^{-4}\Big)\Big]$ (47)
Where we have defined
@@
$\Delta_{T}(\mathbf{x}):=m(\mathbf{x})-\mu_{T}(\mathbf{x}),\quad\Delta_{R}(\mathbf{x}):=m(\mathbf{x})-\mu_{R}(\mathbf{x}),\quad t(\mathbf{x})=\frac{T}{2s(\mathbf{x})^{2}}$ (49)
and all other quantities are as in Result 1.
@@
$f(x)=\log\bar{x}+\frac{1}{\bar{x}}(x-\bar{x})-\frac{1}{2\bar{x}^{2}}(x-\bar{x})^{2}+\frac{1}{3\bar{x}^{3}}(x-\bar{x})^{3}+\mathcal{O}\bigg(\frac{(x-\bar{x})^{4}}{\bar{x}^{4}}\bigg)$ (50)
@@
$\mathbb{E}(\log x)=\log\mathbb{E}(x)-\frac{\mathbb{E}(x^{2})-\mathbb{E}(x)^{2}}{2\mathbb{E}(x)^{2}}+\frac{\mathbb{E}(x^{3})-3\mathbb{E}(x^{2})\mathbb{E}(x)+2\mathbb{E}(x)^{3}}{3\mathbb{E}(x)^{3}}+\mathcal{O}\bigg(\frac{\mathbb{E}((x-\mathbb{E}(x))^{4})}{\mathbb{E}(x)^{4}}\bigg)$ (51)
@@
$\hskip 113.81102pt+\frac{\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p})^{3})-3\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p})^{2})\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p})+2\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p}))^{3}}{3\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p}))^{3}}\bigg)\bigg|_{\mathbf{J}=0}$
@@
$h\!\left(m_{1},m_{2},m_{3};\,s_{1}^{2},s_{2}^{2},s_{3}^{2}\right)=\frac{\exp\!\Bigl[-\dfrac{\displaystyle\sum_{1\leq i\neq j\neq k\leq 3}\bigl(\tfrac{1}{2}m_{i}^{2}(s_{j}^{2}+s_{k}^{2})-m_{i}m_{j}s_{k}^{2}\bigr)}{2\,\prod_{i=1}^{3}s_{i}^{2}\,\displaystyle\sum_{i=1}^{3}\frac{1}{s_{i}^{2}}}\Bigr]}{2\pi\sqrt{\prod_{i=1}^{3}s_{i}^{2}\,\displaystyle\sum_{i=1}^{3}\frac{1}{s_{i}^{2}}}}.$ (56)
@@
$\mathbb{E}_{\mathbf{p}}\!\bigl[z(\mathbf{J},\mathbf{p})\bigr]$ $=\frac{1}{k}\sum_{i=1}^{k}\pi\sqrt{\frac{T}{J_{i}}}\,h\!\Bigl(m,\,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\,\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{i}}\Bigr),$ (57)
$\mathbb{E}_{\mathbf{p}}\!\bigl[z(\mathbf{J},\mathbf{p})^{2}\bigr]$ $=\frac{1}{k^{2}}\Biggl[\sum_{\begin{subarray}{c}i,j=1\\[2.0pt] i\neq j\end{subarray}}^{k}\pi\sqrt{\frac{T}{J_{i}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{i}}\Bigr)\pi\sqrt{\frac{T}{J_{j}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{j}}\Bigr)$
$\hskip 56.9055pt+\sum_{i=1}^{k}\pi\sqrt{\frac{T}{4J_{i}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{4},\tfrac{1}{4J_{i}}\Bigr)\Biggr],$ (58)
@@
$\hskip 66.86414pt\times\pi\sqrt{\frac{T}{J_{j}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{j}}\Bigr)\pi\sqrt{\frac{T}{J_{\ell}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{\ell}}\Bigr)$
$\quad+\sum_{\begin{subarray}{c}i,j=1\\[2.0pt] i\neq j\end{subarray}}^{k}3\,\pi\sqrt{\frac{T}{4J_{i}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{4},\tfrac{1}{4J_{i}}\Bigr)\pi\sqrt{\frac{T}{J_{j}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{j}}\Bigr)$
$\quad+\sum_{i=1}^{k}\pi\sqrt{\frac{T}{9J_{i}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{6},\tfrac{1}{6J_{i}}\Bigr)\Biggr].$ (59)
@@
$\delta(\mathbf{x})$ $=(m-\mathbf{w}_{T}\cdot\frac{\mathbf{x}}{\sqrt{d}})^{2}+s^{2}-\frac{(1-\frac{1}{k})\left((m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}})\left(2(m-\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}})\right)+s^{2}\right)}{t}$ (60)
$+\frac{(1-\frac{1}{k})(1-\frac{2}{k})\left(\left(m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}}\right)\left(2\left(m-\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}}\right)+\left(m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}}\right)\right)+s^{2}\right)}{t^{2}}+O\left(\frac{1}{t^{3}}\right)$
@@
$\delta(\mathbf{x})\;=\;(m-\mathbf{w}_{T}\cdot\frac{\mathbf{x}}{\sqrt{d}})^{2}+s^{2}+\Sigma_{l=1}^{3}(-1)^{l}\Big(1-\tfrac{1}{k}\Big)\Big(1-\tfrac{2}{k}\Big)\dots\Big(1-\tfrac{l}{k}\Big)\frac{1}{t^{l}}C_{l}(\mathbf{x})+O\left(\frac{1}{t^{4}}\right)$ (61)
$\hskip 28.45274ptC_{l}(\mathbf{x})=2\,(m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}})(m-\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}})+s^{2}(\mathbf{x})+(l-1)(m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}})^{2}$ (62)

### v2

@@ A.1 Limit Laws for Maxima
$\mathbb{P}\!\left(\frac{M_{n}-b_{n}}{a_{n}}\leq x\right)\xrightarrow[n\to\infty]{}H(x)$
@@ A.2 Maximum Domains of Attraction and Norming Constants
$\lim_{n\to\infty}F^{n}(a_{n}x+b_{n})=H(x)$
@@
$\lim_{n\to\infty}n\bigl(1-F(a_{n}x+b_{n})\bigr)=-\ln H(x)$
@@ Appendix B Proof of Result 1
$\delta=\mathbb{E}_{\mathbf{x}}\,\mathbb{E}_{y_{1},y_{2},..}\left[\frac{\sum_{i=1}^{k}\big(y_{i}-\mu_{T}(\mathbf{x})\big)^{2}\,e^{-\big(y_{i}-\mu_{R}(\mathbf{x})\big)^{2}/T}}{\sum_{j=1}^{k}e^{-\big(y_{j}-\mu_{R}(\mathbf{x})\big)^{2}/T}}\right],$ (30)
to the leading order in $\sigma$. Here the expectation is over $\mathbf{x}\sim\mathcal{N}(0,\bm{\Sigma}),y_{i}\sim\mathcal{N}(m(\mathbf{x}),s(\mathbf{x})^{2}),i=1,2,\dots,k$. The posterior predictive has mean $m(\mathbf{x})$ and variance $\Sigma(\mathbf{x})=s(\mathbf{x})^{2}$ as follows
@@
$\hat{R}=R\Big(1-\alpha\,m_{\bm{\Sigma}}(R)\Big)=\frac{\sigma^{2}}{\gamma^{2}}\alpha,$ (33)
$m_{\bm{\Sigma}}(R)=\frac{1}{d}\text{Tr}\!\big[\bm{\Sigma}(\bm{\Sigma}+RI)^{-1}\big].$
@@
$m(\mathbf{x})$ $=\bm{\mu}^{\top}\frac{\mathbf{x}}{\sqrt{d}}=\frac{1}{\sigma^{2}}\,\frac{\mathbf{x}}{\sqrt{d}}^{\top}\bm{\Omega}\sum_{i=1}^{n}y^{i}\frac{\mathbf{x}^{i}}{\sqrt{d}}$ (35)
@@
$\mathrm{Var}\big(Z(\mathbf{x})\big)\;\xrightarrow[d,n\to\infty]{}\;\sigma^{2}\,\frac{\alpha\,m_{\bm{\Sigma}}^{(2)}(R)}{1-\alpha\,m_{\bm{\Sigma}}^{(2)}(R)},$ (42)
$m_{\bm{\Sigma}}^{(2)}(R)=\frac{1}{d}\text{Tr}\!\big[\bm{\Sigma}^{2}(\bm{\Sigma}+R\mathbf{I})^{-2}\big]$
@@
$\sum_{i=1}^{n}\frac{\mathbf{x}^{i}}{\sqrt{d}}{\frac{\mathbf{x}^{i}}{\sqrt{d}}}^{\top}=\frac{1}{d}\sum_{i=1}^{n}\mathbf{x}^{i}{\mathbf{x}^{i}}^{\top}=\frac{n}{d}\,\widehat{\bm{\Sigma}}=\frac{1}{\alpha}\,\widehat{\bm{\Sigma}}$, (44)
$\hskip 56.9055pt\bm{\Omega}=\sigma^{2}\alpha(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}.$
@@
$\mathrm{Var}\big(Z(\mathbf{x})\mid\mathcal{D},\mathbf{x}\big)$ (45)
$=\frac{1}{\sigma^{2}\alpha}\,\frac{\mathbf{x}}{\sqrt{d}}^{\top}\bigl[\sigma^{2}\alpha(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\bigr]\widehat{\bm{\Sigma}}\bigl[\sigma^{2}\alpha(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\bigr]\frac{\mathbf{x}}{\sqrt{d}}$
@@
$\mathbb{E}_{\mathbf{x}}\big[\mathrm{Var}(Z(\mathbf{x})\mid\mathcal{D},\mathbf{x})\big]$ (47)
$=\sigma^{2}\alpha\,\frac{1}{d}\,\text{Tr}\Big((\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\bm{\Sigma}\Big).$
@@
$\mathrm{Var}\big(Z(\mathbf{x})\big)$ $=\mathbb{E}_{\mathcal{D},\mathbf{x}}\big[\mathrm{Var}(Z(\mathbf{x})\mid\mathcal{D},\mathbf{x})\big],$ (48)
@@
$\mathrm{Var}\big(Z(\mathbf{x})\big)$ $=\sigma^{2}\alpha\,\frac{1}{d}\,\text{Tr}\Big((\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\widehat{\bm{\Sigma}}(\widehat{\bm{\Sigma}}+\hat{R}\mathbf{I})^{-1}\bm{\Sigma}\Big)$ (51)
$=\sigma^{2}\,\frac{\alpha\,m_{\bm{\Sigma}}^{(2)}(R)}{1-\alpha\,m_{\bm{\Sigma}}^{(2)}(R)}.$
@@ Appendix C Proof of Result 2
$\delta$ $=\;\mathbb{E}_{\mathbf{x}}\Big[\Delta_{T}(\mathbf{x})^{2}+s^{2}(\mathbf{x})$ (53)
$+\Sigma_{l=1}^{3}(-1)^{l}\frac{C_{l}(\mathbf{x})}{t(\mathbf{x})^{l}}\prod_{i=1}^{l}\Big(1-\tfrac{i}{k}\Big)+\mathcal{O}\!\Big(t(\mathbf{x})^{-4}\Big)\Big]$
Where we have defined $t(\mathbf{x})=\frac{T}{2s(\mathbf{x})^{2}}$ and
@@
$\Delta_{T}(\mathbf{x}):=m(\mathbf{x})-\mu_{T}(\mathbf{x}),\quad\Delta_{R}(\mathbf{x}):=m(\mathbf{x})-\mu_{R}(\mathbf{x}).$ (55)
All other quantities are as in Result 1.
@@
$f(x)$ $=\log\bar{x}+\frac{1}{\bar{x}}(x-\bar{x})-\frac{1}{2\bar{x}^{2}}(x-\bar{x})^{2}$ (56)
$+\frac{1}{3\bar{x}^{3}}(x-\bar{x})^{3}+\mathcal{O}\bigg(\frac{(x-\bar{x})^{4}}{\bar{x}^{4}}\bigg)$
@@
$\mathbb{E}(\log x)$ $=\log\mathbb{E}(x)-\frac{\mathbb{E}(x^{2})-\mathbb{E}(x)^{2}}{2\mathbb{E}(x)^{2}}$ (57)
$+\frac{\mathbb{E}(x^{3})-3\mathbb{E}(x^{2})\mathbb{E}(x)+2\mathbb{E}(x)^{3}}{3\mathbb{E}(x)^{3}}+\mathcal{O}\bigg(\frac{\mathbb{E}((x-\mathbb{E}(x))^{4})}{\mathbb{E}(x)^{4}}\bigg)$
@@
$+\frac{\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p})^{3})-3\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p})^{2})\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p})+2\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p}))^{3}}{3\mathbb{E}_{\mathbf{p}}(z(\mathbf{J},\mathbf{p}))^{3}}\bigg)\bigg|_{\mathbf{J}=0}$
@@
$h\!\left(m_{1},m_{2},m_{3};\,s_{1}^{2},s_{2}^{2},s_{3}^{2}\right)$ (62)
$=\frac{\exp\!\Bigl[-\dfrac{\displaystyle\sum_{1\leq i\neq j\neq k\leq 3}\bigl(\tfrac{1}{2}m_{i}^{2}(s_{j}^{2}+s_{k}^{2})-m_{i}m_{j}s_{k}^{2}\bigr)}{2\,\prod_{i=1}^{3}s_{i}^{2}\,\displaystyle\sum_{i=1}^{3}\frac{1}{s_{i}^{2}}}\Bigr]}{2\pi\sqrt{\prod_{i=1}^{3}s_{i}^{2}\,\displaystyle\sum_{i=1}^{3}\frac{1}{s_{i}^{2}}}}.$
@@
$\mathbb{E}_{\mathbf{p}}\!\bigl[z(\mathbf{J},\mathbf{p})\bigr]$ (63)
$=\frac{1}{k}\sum_{i=1}^{k}\pi\sqrt{\frac{T}{J_{i}}}\,h\!\Bigl(m,\,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\,\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{i}}\Bigr),$
$\mathbb{E}_{\mathbf{p}}\!\bigl[z(\mathbf{J},\mathbf{p})^{2}\bigr]$ $=\frac{1}{k^{2}}\Biggl[\sum_{\begin{subarray}{c}i,j=1\\[2.0pt] i\neq j\end{subarray}}^{k}\pi\sqrt{\frac{T}{J_{i}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{i}}\Bigr)$
$\hskip 28.45274pt\times\pi\sqrt{\frac{T}{J_{j}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{j}}\Bigr)$
$+\sum_{i=1}^{k}\pi\sqrt{\frac{T}{4J_{i}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{4},\tfrac{1}{4J_{i}}\Bigr)\Biggr],$
@@
$\hskip 28.45274pt\times\pi\sqrt{\frac{T}{J_{j}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{j}}\Bigr)$
$\hskip 28.45274pt\times\pi\sqrt{\frac{T}{J_{\ell}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{\ell}}\Bigr)$
$\quad+\sum_{\begin{subarray}{c}i,j=1\\[2.0pt] i\neq j\end{subarray}}^{k}3\,\pi\sqrt{\frac{T}{4J_{i}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{4},\tfrac{1}{4J_{i}}\Bigr)$
$\hskip 28.45274pt\times\pi\sqrt{\frac{T}{J_{j}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{2},\tfrac{1}{2J_{j}}\Bigr)$
$\quad+\sum_{i=1}^{k}\pi\sqrt{\frac{T}{9J_{i}}}\,h\!\Bigl(m,\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}},\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}};s^{2},\tfrac{T}{6},\tfrac{1}{6J_{i}}\Bigr)\Biggr].$
@@
$\delta(\mathbf{x})$ $=(m-\mathbf{w}_{T}\cdot\frac{\mathbf{x}}{\sqrt{d}})^{2}+s^{2}$ (64)
$-\frac{1}{t}\left(1-\frac{1}{k}\right)\left((m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}})\left(2(m-\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}})\right)+s^{2}\right)$
$+\frac{1}{t^{2}}\left(1-\frac{1}{k}\right)\left(1-\frac{2}{k}\right)$
$\times\left(\left(m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}}\right)\left(2\left(m-\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}}\right)+\left(m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}}\right)\right)+s^{2}\right)$
$\hskip 56.9055pt+O\left(\frac{1}{t^{3}}\right)$
@@
$\delta(\mathbf{x})$ $=\;(m-\mathbf{w}_{T}\cdot\frac{\mathbf{x}}{\sqrt{d}})^{2}+s^{2}$ (65)
$+\Sigma_{l=1}^{3}(-1)^{l}\Big(1-\tfrac{1}{k}\Big)\Big(1-\tfrac{2}{k}\Big)\dots\Big(1-\tfrac{l}{k}\Big)\frac{1}{t^{l}}C_{l}(\mathbf{x})$
$+O\left(\frac{1}{t^{4}}\right)$
$C_{l}(\mathbf{x})$ $=2\,(m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}})(m-\frac{\mathbf{w}_{\!T}\cdot\mathbf{x}}{\sqrt{d}})$
$+s^{2}(\mathbf{x})+(l-1)(m-\frac{\mathbf{w}_{\!R}\cdot\mathbf{x}}{\sqrt{d}})^{2}$